OpenStudy (anonymous):

please help me with x=(-3±√(3^2-4(6)(-18)))/(2(6)) I got as far as x= -3±√441/12 but then get stuck there

OpenStudy (anonymous):

the square root of 441 is 21

OpenStudy (anonymous):

the square root of 441 is 21 so we have x = (-3 +21)/12 = 3/2 and x = (-3 -21)/12 = -2

OpenStudy (anonymous):

you were almost there

OpenStudy (anonymous):

I get confused on the first x = (-3+21)/12 because its a fraction

OpenStudy (anonymous):

I end up with x = 18/12 but then what

OpenStudy (anonymous):

reduce by 6

OpenStudy (anonymous):

-3 + 21 = 18 so we have 18/12 18 and 12 are divisible by 6 18 divided by 6 = 3 12 divided by 6 = 2 so we have 18/12 = 3/2

OpenStudy (anonymous):

oh my gosh, duh...lol i think i think too hard about somethings and miss the obvious

OpenStudy (anonymous):

thank you very much